Machine checkout meeting – Collimator 20.05.2009

sunglowcitrineUrban and Civil

Nov 15, 2013 (3 years and 1 month ago)

61 views

Mach
ine checkout meeting


Collimator

20
.0
5
.2009

Presents: O. Aberle, R. Alemany
-
Fernandez, G. Arduini, R. Assmann, R. Gi
achin
o,
S.

Redaelli


Collimator installation status:

22 collimators have been changed in the ring and 6 in the transfer lines. Installation almost finished
and database
being

updated. The only technical issue found so far is relate
d to problems with some
springs that pre
-
load the jaws. This leads to mechanic
al play (up to 130 micrometers). Some tests
will be done next week and actions will be taken if needed.

Collimation commissioning;

Individual system test:



Middle of June.

Although the hardware commissioning won’t be fully finished, the collimators will be

tested
from the CCC during a week.
Participants
: collimation team.



2
nd

week of August (w33)

Collimation system will be fr
ozen

and final commissioning in REMOTE will take place to be
prepared for operation with beam.
Participants
: collimation team + OP.

Ma
nagement of critical settings (MCS):

MCS tests have been performed with the expert tools which don’t go through TRIM. A test going
through TRIM has to be scheduled to validate the final procedure. For the moment MCS is not used
otherwise the collimator exp
ert tools cannot be used.

Action: R
alph

Machine Protection Commissioning procedures need to be finalized.
Action: Ralph

AUG tests have been performed but it is not clear whether the collimator experts have been
involved.
Action: Gianluigi to verify with Hu
gues.

The movement of the TOTEM Roman pots is exactly the same as that of the col
lima
tors. Interlock
tests will be done as part of the collimator tests.

Action: Ralph.

What about ATLAS?

to be verified by Gianluigi with Ralph.

Status of the Alarms (e
-
mail f
rom Markus):

1) The alarm category "LHC Collimators" has been defined below the node "LHC
Operation"

2) The above category is populated with all motorisation alarms coming from
the respective FESA class.


Outstanding issues:


1) The alarm priority for the motorisation alarms is currently set to 2 which
is mapped to yellow for the LHC. The collimator experts should decide for
which of the alarms this needs to be changed


Action: Markus,
Stefano

2) The temperature alarms are not

yet implemented. Enrique Blanco is in charge
of this and will inform u
s as soon as they are available


Action: Markus,
Enrique

3) No reduction alarms have been set up so far. This will be done if necessary
after consultation with the collimator experts.

4) Alarm tes
ts need to be organised, prefer
ably during a collimator dry run.

Action: Ralph, Markus


The list of logged parameters is available. Is it the last one
?
Are the post
-
mortem data defined?

Action: Gianluigi and Stefano to verify.

It is not yet cle
ar who is going to provide the tools for the alignment of the collimator jaws based on
BLM measurements.

Action: Ralph.

A first version of the fixed display exists. This provides the status of the
collimators and their
position.
Possible other parameters
to be added: BLM signals, temperature of the jaws, vacuum,
experiment background, beam intensity and/or lifetime.

TDI and TDCQ: check of the collimator interlocks (temperature and position) will be done by the
collimation team the machine protection functi
onality will be tested by ABT.

A.O.B.

After June another machine checkout meeting will have to be organized to finalize the procedures
and see what will be left for the full machine checkout period.

R. Alemany, 04
/
06
/2009